OpenStudy (anonymous):

In the function f(x) = 5x + 12 how does an increase of 1 in the independent variable x affect f(x)? f(x) will have an increase of 5 f(x) will have a decrease of 17 f(x) will have an increase of 5x + 5 f(x) will have a decrease of 5

OpenStudy (anonymous):

increase by 5

OpenStudy (anonymous):

if x = 2 you get \[5\times 2+12=10+12=22\] if x = 3 you get \[5\times 3+12=15+12=27\]

OpenStudy (anonymous):

x increases by 1, y increases by 5

OpenStudy (anonymous):

thanks:))

OpenStudy (anonymous):

yw